# Apparent Selection Pressure for Channel Capacity in Bacterial Chemotactic Sensors

### v2

@@ IV.1 Channel capacity
For all strains, the capacity-achieving input distribution $p^{*}(c)$ is strongly bi-modal, concentrating probability near the two endpoints of the response and showing a pronounced minimum at the midpoint. See Fig. 4. For a noisy binary receptor, observations are most informative when the output is near-deterministic ($p_{\mathrm{active}}(c)\approx 0$ or $\approx 1$) and least informative at the midpoint where the Bernoulli variance p(1p)p(1-p) is maximal. In other words, the information-maximizing strategy is spend time in both low and high concentration regions rather than to dwell at the middle. The corresponding optimal output distribution $p^{*}(s)$ is slightly biased toward the active state, which indicate mild channel asymmetry.
@@
All optimal input distributions are similar despite differences in strain parameters. Figure 5 thus forms a testable prediction for these strains. Actual input distributions in natural environments in which gradients constantly vanish [4] might resemble the optimal input distributions shown here, with mutual informations close to capacity.
@@
Figure 4: Optimal use of the bacterial chemotactic receptors for all strains using parameters in Table 1. At top, capacity-achieving input distribution $p^{*}(c)$ versus ligand concentration $c$ as obtained by the Blahut-Arimoto algorithm. At bottom, optimal output distribution $p^{*}(s)$ among the active and inactive state for all strains.
@@
Figure 5: Channel capacity versus $K_{d}^{(A),1}$ for each of the ten strains holding all other parameters fixed.
Strain $C$ for strain $_{max} CC$ for strain / $C_{max}$
Strain 1 0.83 0.95 0.87
Strain 2 0.76 0.95 0.80
Strain 3 0.64 0.95 0.68
Strain 4 0.81 0.95 0.86
Strain 5 0.65 0.95 0.68
Strain 6 0.80 0.95 0.85
Strain 7 0.87 0.95 0.92
Strain 8 0.77 0.95 0.81
Strain 9 0.82 0.95 0.86
Strain 10 0.86 0.95 0.90
@@
Table 3: Channel capacity by strain. With $C_{max}$ the maximum channel capacity found over the 7D parameter sweep, the channel capacity for each strain. All capacities are in bits.
Strain Gradient $\|\nabla C\|_{2}$ for strain Global max $\|\nabla C\|_{2}$ Gradient norm ratio
Strain 1 0.02 5.75 0.00
Strain 2 0.10 5.75 0.02
Strain 3 0.47 5.75 0.08
Strain 4 0.06 5.75 0.01
Strain 5 0.62 5.75 0.11
Strain 6 0.12 5.75 0.02
Strain 7 0.14 5.75 0.02
Strain 8 0.40 5.75 0.07
Strain 9 0.15 5.75 0.03
Strain 10 0.11 5.75 0.02
@@
Table 4: Local flatness of channel capacity as measured by the L2 norm of the gradient of channel capacity. Gradient L2 norm for each strain versus the global maximum of the gradient L2 norm over the seven-dimensional parameter sweep. All gradients estimated numerically.
@@
Figure 5 maps channel capacity $C$ across a key MWC parameter. A binary receptor conveys information only if there exist input concentrations whose outputs are meaningfully different, which is determined by the dynamic range. When the two endpoints are well separated, the Blahut–Arimoto algorithm concentrates input probability near those regions, making the outputs easy to distinguish and yielding a high $C$; when the dynamic range is small, all inputs appear similar at the output and $C$ collapses. Because channel capacity is strongly correlated with dynamic range ($R=0.978$), the dominant qualitative parameter dependencies of $C$ can be understood directly from the dynamic range expression, and we therefore defer a detailed discussion of these trends to the following subsection.
Fig. 5 suggests that strains are poised at a local maximum of channel capacity in seven-dimensional parameter space. This is confirmed by numerical calculation of gradient norms of channel capacity versus the maximum possible gradient norm of channel capacity in Table 4. However, we are at a local maximum rather than a global maximum of channel capacity based on Table 3.
@@ IV.2 Dynamic range
Since dynamic range and channel capacity are strongly correlated (Table 2), channel capacity can be observed following the same qualitative trends.
Dynamic range looks, based on representative heatmaps, to be maximized in parameter space in Fig. 6. Like for channel capacity, it looks like strains sit on the edge of a plateau for dynamic range. And according to Table 5, this dynamic range maximization is global, with most strain’s dynamic ranges near the global maximum of $1$. However, Table 6 tells a slightly different story: Numerical gradients of dynamic range are very large for many of the strains, so the heatmaps are actually misleading upon closer inspection. From this, we conclude that dynamic range is large and but potentially not maximized for these strains.
Figure 6: A representative heatmap suggesting that dynamic range is also at a global maximum in parameter space, though see Table 6. Dynamic range versus $K_{d}^{(A),1}$ and $K_{d}^{(A),2}$ appears to be locally maximized by strain parameters for strain 11.
Strain $\mathrm{DR}$ for strain $\mathrm{DR}_{max}$ $\mathrm{DR}$ for strain / $\mathrm{DR}_{max}$
@@
Table 5: Dynamic range by strain compared with sweep maxima. $\mathrm{DR}_{max}$ is the maximum dynamic range over the 7D sweep.
Strain Gradient $\|\nabla\mathrm{DR}\|_{2}$ for strain Global max of $\|\nabla\mathrm{DR}\|_{2}$ (7D) Gradient norm ratio
Strain 1 0.27 6.00 0.05
Strain 2 0.29 6.00 0.05
Strain 3 0.47 6.00 0.08
Strain 4 0.69 6.00 0.12
Strain 5 2.71 6.00 0.45
Strain 6 2.98 6.00 0.50
Strain 7 3.00 6.00 0.50
Strain 8 1.62 6.00 0.27
Strain 9 1.87 6.00 0.31
Strain 10 2.33 6.00 0.39
Table 6: Lack of flatness of dynamic range as measured by the L2 norm of the gradient of dynamic range. Gradient L2 norm for each strain versus the global maximum of the gradient L2 norm over the seven-dimensional parameter sweep. All gradients estimated numerically.
@@ IV.3 Effective Hill coefficient
Based on Fig. 7, it is not locally maximized for strain 1; and see Supplementary Information for other strains. This intuition is borne out by Table 7, which shows that effective Hill coefficients are far from a global maximum, and Table 8, which shows that strains have large Hill coefficient gradient norms. From this, we conclude that effective Hill coefficient is not locally or globally maximized.
Figure 7: A heatmap suggesting that effective Hill coefficient is not at a local maximum in parameter space. Effective Hill coefficient versus $K_{d}^{(A),1}$ and $K_{d}^{(A),2}$ is locally maximized by strain parameters for strain 11.
Strain $n_{\mathrm{eff}}$ at bio dot Max $n_{\mathrm{eff}}$ over 7D sweep Ratio of $n_{\mathrm{eff}}$
Strain 1 5.03 176.96 0.03
Strain 2 2.74 176.96 0.02
Strain 3 1.87 176.96 0.01
Strain 4 4.48 176.96 0.03
Strain 5 2.23 176.96 0.01
Strain 6 4.46 176.96 0.03
Strain 7 11.10 176.96 0.06
Strain 8 2.63 176.96 0.01
Strain 9 3.94 176.96 0.02
Strain 10 8.43 176.96 0.05
Table 7: Effective Hill coefficient by strain compared with sweep maxima.
Strain Gradient $\|\nabla n_{\mathrm{eff}}\|_{2}$ at bio point Global max $\|\nabla n_{\mathrm{eff}}\|_{2}$ (7D) Ratio of gradient norms
Strain 1 9.54 903.03 0.01
Strain 2 3.83 903.03 0.00
Strain 3 2.29 903.03 0.00
Strain 4 9.10 903.03 0.01
Strain 5 3.86 903.03 0.00
Strain 6 10.58 903.03 0.01
Strain 7 30.48 903.03 0.03
Strain 8 4.53 903.03 0.01
Strain 9 8.29 903.03 0.01
Strain 10 21.81 903.03 0.02
Table 8: Local flatness of the effective Hill coefficient: gradient norm at each strain versus the global maximum over the 7D sweep.
@@ V Conclusions
This study examined the static sensing limits of mixed Tar/Tsr chemoreceptor clusters in E. coli using a heterogeneous MWC framework. By sweeping a seven-parameter space and computing three performance metrics—channel capacity, effective Hill coefficient, and dynamic range—we mapped how receptor composition and state-dependent affinities shape the information available at the receptor output level. While dynamic range tends to be large, it does not seem to be maximized based on gradient analyses, and the effective Hill coefficient appears to not be maximized either. Channel capacity reveals itself as locally maximized by all E. coli strains. This is revealed even though all three characterizations of performance are correlated, as in Ref. [9].

### v3

@@ IV.1 Channel capacity
For all strains, the capacity-achieving input distribution $p_{\mathrm{in}}^{*}(c)$ is strongly bimodal, concentrating probability near the two endpoints of the response and showing a pronounced minimum at the midpoint (Fig. 4). For a noisy binary receptor, observations are most informative when the output is near-deterministic ($p(c)\approx 0$ or $\approx 1$) and least informative at the midpoint where the Bernoulli variance p(1p)p(1-p) is maximal. In other words, the information-maximizing strategy is to spend time in both low and high concentration regions rather than to dwell in the middle. The corresponding optimal output distribution $p^{*}(s)$ is slightly biased toward the active state, which indicates mild channel asymmetry.
@@
All optimal input distributions are similar despite differences in strain parameters. Fig. 4 thus forms a testable prediction for these strains. The actual input distributions in natural environments in which gradients constantly vanish [14] might resemble the optimal input distributions shown here, with mutual information values close to capacity.
@@
Figure 4: Optimal use of the bacterial chemotactic receptors for all strains using parameters in Table 1. At top, capacity-achieving input distribution $p_{\mathrm{in}}^{*}(c)$ versus ligand concentration $c$ as obtained by the Blahut-Arimoto algorithm. At bottom, optimal output distribution $p^{*}(s)$ among the active and inactive state for all strains.
@@
Figure 5: Channel capacity versus $K_{d}^{(A),1}$ for each of the ten strains while holding all other parameters fixed.
Strain $C_{\text{strain}}$ $C_{\text{max}}$ $C_{\text{strain}}$ / $C_{\text{max}}$
Strain 1 0.99 1.00 0.99
Strain 2 0.93 1.00 0.93
Strain 3 0.80 1.00 0.80
Strain 4 0.99 1.00 0.99
Strain 5 0.83 1.00 0.83
Strain 6 0.98 1.00 0.98
Strain 7 1.00 1.00 1.00
Strain 8 0.93 1.00 0.93
Strain 9 0.99 1.00 0.99
Strain 10 1.00 1.00 1.00
@@
Table 3: Channel capacity $C_{\text{strain}}$ for each strain operating point, compared with the maximum $C_{\max}$ over the 7D parameter sweep. Capacities are in bits.
Strain $\|\nabla C\|_{2}$ $\|\nabla\mathrm{DR}\|_{2}$ $\|\nabla|n_{\mathrm{eff}}|\|_{2}$
Strain 1 0.01 0.00 6.10
Strain 2 0.12 0.04 3.01
Strain 3 0.34 0.15 2.04
Strain 4 0.02 0.00 5.17
Strain 5 0.27 0.13 3.03
Strain 6 0.02 0.01 5.93
Strain 7 0.00 0.00 14.90
Strain 8 0.15 0.04 3.81
Strain 9 0.03 0.00 5.94
Strain 10 0.00 0.00 12.94
@@
Table 4: Local flatness of channel capacity, dynamic range, and effective Hill coefficient as measured by the $L_{2}$ norm of the numerical gradient of each performance metric at each strain operating point. All gradients are estimated numerically.
@@
Figure 5 maps channel capacity $C$ across a key MWC parameter. A binary receptor conveys information only if there exist input concentrations whose outputs are meaningfully different, which is determined by the dynamic range. When the two endpoints are well separated, the Blahut-Arimoto algorithm concentrates input probability near those regions, making the outputs easy to distinguish and yielding a high $C$; when the dynamic range is small, all inputs appear similar at the output and $C$ collapses. Because channel capacity is strongly correlated with dynamic range ($R=0.990$), the dominant qualitative parameter dependencies of $C$ can be understood directly from the dynamic range expression, and we therefore defer a detailed discussion of these trends to the following subsection.
Fig. 5 indicates that strains are poised near a global maximum of channel capacity in seven-dimensional parameter space, as evidenced by the small gradient norms at the strain operating points (Table 4) and the high channel capacities in Table 3. Taken together, these results indicate that real E. coli strains are tuned to a global optimum.
@@ IV.2 Dynamic range
In Fig. 6, the dynamic-range landscape exhibits a broad near-maximal plateau, with the strain-1 operating point lying near its boundary. Quantitatively, Table 5 shows that most strains achieve dynamic range values close to the global maximum of $1$. Consistent with this, the gradient norm in Table 4 indicates that the local gradient at the strain operating points is small, supporting the interpretation that dynamic range is maximized globally for these bacterial strains.
Figure 6: Representative heatmap of dynamic range $\mathrm{DR}$ versus $K_{d}^{(A),1}$ and $K_{d}^{(A),2}$, suggesting that the strain-1 operating point lies near the global maximum of dynamic range.
Strain $\mathrm{DR}_{\text{strain}}$ $\mathrm{DR}_{\text{max}}$ $\mathrm{DR}_{\text{strain}}$ / $\mathrm{DR}_{\text{max}}$
@@
Table 5: Dynamic range $\mathrm{DR}_{\text{strain}}$ for each strain operating point, compared with the maximum $\mathrm{DR}_{\max}$ over the 7D parameter sweep.
@@ IV.3 Effective Hill coefficient
Based on Fig. 7, effective Hill coefficient is not locally maximized for strain 1. This intuition is borne out by Table 6, which shows that effective Hill coefficients are far from a global maximum, and Table 4, which shows that strains have large Hill coefficient gradient norms. From this, we conclude that effective Hill coefficient is not locally or globally maximized.
Figure 7: Representative heatmap of effective Hill coefficient $n_{\mathrm{eff}}$ versus $K_{d}^{(A),1}$ and $K_{d}^{(A),2}$, suggesting that $n_{\mathrm{eff}}$ is not locally maximized at the strain-1 operating point.
Strain $\lvert n_{\mathrm{eff}}\rvert_{\text{strain}}$ $\lvert n_{\mathrm{eff}}\rvert_{\text{max}}$ $\lvert n_{\mathrm{eff}}\rvert_{\text{strain}}$ / $\lvert n_{\mathrm{eff}}\rvert_{\text{max}}$
Strain 1 5.09 62.82 0.08
Strain 2 2.74 62.82 0.04
Strain 3 1.96 62.82 0.03
Strain 4 4.48 62.82 0.07
Strain 5 2.23 62.82 0.04
Strain 6 4.45 62.82 0.07
Strain 7 11.08 62.82 0.18
Strain 8 2.60 62.82 0.04
Strain 9 3.97 62.82 0.06
Strain 10 8.48 62.82 0.13
Table 6: Effective Hill coefficient $\lvert n_{\mathrm{eff}}\rvert_{\text{strain}}$ for each strain operating point, compared with the maximum $\lvert n_{\mathrm{eff}}\rvert_{\text{max}}$ over the 7D parameter sweep.
@@ V Conclusions
This study examined the static sensing limits of mixed Tar/Tsr chemoreceptor clusters in E. coli using a heterogeneous MWC framework. By sweeping a seven-parameter space and computing three performance metrics—channel capacity, effective Hill coefficient, and dynamic range—we mapped how receptor composition and state-dependent affinities shape the information available at the receptor output level. The effective Hill coefficient does not appear to be maximized. Channel capacity reveals itself as globally maximized, and dynamic range too, across all E. coli strains considered. This is revealed even though all three characterizations of performance are correlated, as in Ref. [3].
@@
Acknowledgements.
Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-25-1-0260. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. All code is available at https://github.com/Hail-Earendil/bacterial-chemotaxis-7d-sweep.